Please briefly describe the physical significance of Faraday’s law of electromagnetic induction.


1
Expert's answer
2021-06-23T13:08:58-0400

states that the induced voltage in a circuit is proportional to the rate of change over time of the magnetic flux through that circuit. In other words, the faster the magnetic field changes, the greater will be the voltage in the circuit




magnetic field B, the area of the loop A, and the normal n to that area, as represented by the fingers of the left hand. If ΔΦB is positive, the direction of the EMF is the same as that of the curved fingers (yellow arrowheads). If ΔΦB is negative, the direction of the EMF


"e=-\frac{d\phi}{dt}"

When n loop peresent

"e=-N\frac{d\phi}{dt}"

"e=-N\frac{d(BA)}{dt}"

Maxwell Faraday equation

"\nabla\times E=-\frac{dB}{dt}"